which product represents the solution to the system: -y+7x=14
-x+4y=1

Answer by checkley77(12844)

-y+7x=14 multiply by 4 & add.
-x+4y=1
28x-4y=56
---------------------
27x=57
x=57/27
x=19/9 ans.
-19/9+4y=1
4y=1+19/9
4y=(9+19)/9
4y=28/9
y=(28/9)/4
y=(28/9)(1/4)
y=28/36
y=7/9 ans.
-7/9+7(19/9)=14
-7/9+133/9=14
126/9=14
14=14